# DisCoRD: Discrete Tokens to Continuous Motion via Rectified Flow Decoding

### v1

@@ 4 Experiments
In this section, we evaluate the effectiveness of DisCoRD in achieving motion naturalness compared to other discrete methods. We begin by assessing the naturalness of reconstructed motions to highlight the expressive capabilities of our score-based decoder. Then, we examine how this naturalness carries over to stage 2, generating natural motions while preserving faithfulness. We focus on text-to-motion generation due to its complex motions and diversity, but also evaluate our approach on other motion generation tasks, including co-speech gesture generation and music-to-dance generation, demonstrating the flexibility of our method.
@@ 4.1 Dataset and Evaluation
Datasets. For the text-to-motion task, we employed HumanML3D [11] and KIT-ML [37]. HumanML3D is a 3D motion dataset with language annotations, including 14,616 motion sequences paired with 44,970 text descriptions averaging 12 words. Sourced from motion capture data, motions are standardized to a skeletal template, scaled to 20 FPS, and cropped to 10 seconds if longer. Each sequence has at least three descriptive texts covering diverse actions. KIT-ML is a smaller dataset with 3,911 motion sequences paired with 6,278 text descriptions averaging 8 words. Motion capture data are downsampled to 12.5 FPS, with 1–4 descriptions per sequence. For co-speech gesture generation, we utilize the SHOW [58] dataset, while a mixed version of AIST++ [20] and HumanML3D is used for music-to-dance generation. Further dataset details are provided in the Supplementary Section B.
@@
We evaluate DisCoRD on both motion reconstruction and motion generation separately. For motion reconstruction, the primary objective is to assess how effectively the decoder reconstructs motion from tokens. This is measured by Fréchet Inception Distance (FID), which assesses motion realism by comparing the feature distributions of generated and ground truth motions, and Mean Per Joint Position Error (MPJPE), which quantifies positional accuracy. For text-to-motion generation, we follow prior works [44] and we employ several established metrics: FID; R-Precision, which measures retrieval accuracy by matching generated motions to corresponding text descriptions and reports Top-1, Top-2, and Top-3 accuracies; Multimodal Distance (MM-Dist), which calculates the average feature distance between generated motions and associated text descriptions, capturing cross-modal alignment; and Multimodality (MModality), which assesses the model’s capacity to produce diverse motions from a single text prompt by averaging distances between multiple generations for the same prompt. For co-speech gesture generation, we employ Fréchet Gesture Distance (FGD) [31], and for music-to-dance generation, following [46], we utilize Distk and Distg to assess the quality of generated motions by comparing the distributional spread of generated and real motions. Additional specifics on these evaluation metrics are provided in the Supplementary Section B.
Figure 3: sJPE and FID response to gaussian noise. Noise sJPE is very sensitive to small level frame-wise noise than FID, while Static sJPE remains small. Note that FID values are very small.
Symmetric Jerk Percentage Error
Motion evaluation at stage 1 often relies on Mean Per Joint Position Error (MPJPE) and Fréchet Inception Distance (FID). However, MPJPE lacks alignment with human perceptual preferences [8], and FID, as a model-level metric derived from pretrained network features, fails to capture sample-wise naturalness effectively [50]. Our findings further reveal that FID is insensitive to subtle, fine-grained noise, as shown in Figure 3—an essential aspect of immersive motion quality [53]. To address these limitations, we propose an evaluation approach targeting both under-reconstructed motions and frame-wise noise by incorporating jerk. Jerk has proven effective for assessing motion quality [4, 7] and serves as a key indicator of kinetic inconsistencies [3, 19], making it a valuable metric for capturing nuanced deviations in motion. It can be obtained by the third derivative of position with respect to time, effectively capturing the rate of change of acceleration, allowing us to quantify subtle deviations and inconsistencies in motion.
Let $J_{\text{pred},t}$ and $J_{\text{true},t}$ denote the predicted and ground truth jerk, respectively, at time $t$ over $n$ time points. Then the symmetric Jerk Percentage Error (sJPE), capturing the average deviation between predicted and ground truth jerk values, is defined as
@@
Within sJPE, we identify two components: Noise sJPE and Static sJPE. Noise sJPE corresponds to instances where $J_{\text{pred},t}>J_{\text{true},t}$, indicating an overestimation of jerk in the predicted motion, which reflects the presence of fine-grained noise. This effect is evident in discrete-based methods, where discrete tokens introduce frame-wise noise, as shown in Figure 4. Static sJPE captures instances where $J_{\text{pred},t}\leq J_{\text{true},t}$, indicating underestimation of jerk, or insufficiently dynamic predicted motion, highlighted by green boxes in Figure 4. Together, these components provide a comprehensive measure of prediction accuracy, capturing both over- and underestimations of jerk within a unified score. See Supplementary Section C for additional details on sJPE and qaulitatve results.
Figure 4: Under-reconstruction and frame-wise noise. We visualize joint positions (top) and corresponding jerk graphs (bottom). Compared to MoMask, DisCoRD demonstrates lower noise levels (Noise sJPE, blue area) and reduced under-reconstruction (Static sJPE, red area), with under-reconstruction regions in Momask highlighted in green boxes. This results in a lower sJPE for our method, indicating improved naturalness in the reconstructed motion.
@@ 4.2 Quantitative results
HumanML3D
MLD (cont.) 0.017 14.7 0.404
T2M-GPT ${0.089}^{\pm{.001}}$ 60.0 0.564
+DisCoRD(Ours) $\mathbf{{0.031}}^{\pm{.001}}$ 71.5 0.488
MMM ${0.097}^{\pm{.001}}$ 46.9 0.517
+DisCoRD(Ours) $\mathbf{{0.020}}^{\pm{.001}}$ 56.8 0.429
MoMask ${0.019}^{\pm{001}}$ 29.5 0.512
+DisCoRD(Ours) $\mathbf{{0.011}}^{\pm{000}}$ 33.3 0.385
KIT-ML
T2M-GPT ${0.470}^{\pm{.010}}$ 46.4 0.526
+DisCoRD(Ours) $\mathbf{{0.284}}^{\pm{.009}}$ 58.7 0.395
MoMask ${0.113}^{\pm{.001}}$ 37.5 0.384
+DisCoRD(Ours) ${\textbf{0.103}}^{\pm{.002}}$ 33.0 0.359
Table 1: Quantitative evaluation on the HumanML3D and KIT-ML test sets for reconstruction. DisCoRD serves as a discrete model’s decoder, outperforming MLD in both FID and sJPE, demonstrating its effectiveness in decoding discrete tokens in continuous space.
Datasets Methods R Precision $\uparrow$ FID $\downarrow$ MultiModal Dist $\downarrow$ MultiModality $\uparrow$
@@
+ DisCoRD (Ours) ${0.476}^{\pm{.008}}$ ${0.663}^{\pm{.006}}$ ${0.760}^{\pm{.007}}$ ${0.095}^{\pm{.011}}$ ${3.121}^{\pm{.009}}$ ${1.831}^{\pm{.048}}$
MoMask [13] $\underline{{0.521}}^{\pm{.002}}$ $\underline{{0.713}}^{\pm{.002}}$ $\underline{{0.807}}^{\pm{.002}}$ $\underline{{0.045}}^{\pm{.002}}$ $\underline{{2.958}}^{\pm{.008}}$ ${1.241}^{\pm{.040}}$
+ DisCoRD (Ours) $\mathbf{{0.524}}^{\pm{.003}}$ $\mathbf{{0.715}}^{\pm{.003}}$ $\mathbf{{0.809}}^{\pm{.002}}$ $\mathbf{{0.032}}^{\pm{.002}}$ $\mathbf{{2.938}}^{\pm{.010}}$ ${1.288}^{\pm{.043}}$
@@
+ DisCoRD (Ours) ${0.382}^{\pm{.007}}$ ${0.590}^{\pm{.007}}$ ${0.715}^{\pm{.004}}$ ${0.541}^{\pm{.038}}$ ${3.260}^{\pm{.028}}$ $\underline{{1.928}}^{\pm{.059}}$
@@
+ DisCoRD (Ours) $\mathbf{{0.434}}^{\pm{.007}}$ $\mathbf{{0.657}}^{\pm{.005}}$ $\underline{{0.775}}^{\pm{.004}}$ $\underline{{0.169}}^{\pm{.010}}$ $\underline{{2.792}}^{\pm{.015}}$ ${1.266}^{\pm{.046}}$
Table 2: Quantitative evaluation on the HumanML3D and KIT-ML test set. $\pm$ indicates a 95% confidence interval. +DisCoRD indicates that the baseline model’s decoder is replaced with DisCoRD. Bold indicates the best result, while underscore refers the second best. DisCoRD improves naturalness, as evidenced by FID scores, while preserving faithfulness, demonstrated by R-Precision and MM Distance.
@@
To assess the expressiveness of our decoding strategy, we replaced the decoders of state-of-the-art text-to-motion discrete methods with DisCoRD and evaluated their performance in stage 1. Shown in Table 1, while MLD benefits from a continuous representation that enhances naturalness, discrete methods often struggle to reconstruct smooth and natural motions, resulting in increased sJPE values. Although models like MoMask achieve competitive FID, they fall short of MLD in motion naturalness, as indicated by significantly higher sJPE. Our approach yields substantial improvements in both FID and sJPE when applied to discrete models and even surpasses MLD in sJPE, demonstrating that leveraging discrete tokens to decode motion in continuous space combines the advantages of both methods effectively.
@@
Table 3: Quantitative results on co-speech gesture generation. DisCoRD outperforms baseline models on sJPE and FGD.
@@
Table 4: Quantitative results on music-to-dance generation. DisCoRD outperforms baseline model on sJPE, $\mathrm{Dist_{k}}$ and $\mathrm{Dist_{g}}$.
@@
To evaluate DisCoRD’s effectiveness in decoding predicted tokens, we leveraged models trained in stage 1 to assess their performance in decoding tokens generated by pretrained token predictors. As shown in Table 2, our method outperforms baseline models, particularly in FID, achieving state-of-the-art performance and demonstrating that an expressive decoder enhances both reconstruction quality and motion generation. While R-precision scores indicate a decrease in faithfulness for T2M-GPT, our results with MoMask—paired with a powerful tokenizer, RVQVAE—show that DisCoRD attains high naturalness without sacrificing faithfulness, highlighting its potential as a default decoder replacement for discrete methods.
@@
To validate our approach as a general method for enhancing naturalness in discrete-based human motion generation, we trained DisCoRD on co-speech gesture and music-driven dance generation, conducting a comparative analysis against baseline models. As shown in Tables 3 and 4, our method consistently outperforms baseline models across both tasks, achieving superior performance on sJPE and standard evaluation metrics. We present additional evaluation results in Supplementary Section D.
@@
By selecting the rectified flow algorithm for our score-based model, we exploit its efficient transport mechanism to achieve inference speeds comparable to baseline models. As shown in Figure 5, we evaluated the decoding times for both Momask and our model on tokens generated by a pretrained token generator. At the default setting of 16 sampling steps, our model achieves decoding speeds on par with MoMask while delivering superior sJPE, stage 1 FID, and stage 2 FID. Furthermore, by reducing the sampling steps, our method can decode tokens significantly fasterer than MoMask, maintaining comparable or enhanced FID and sJPE performance. Additionally, although it was not the primary focus of this experiment, we observed that sJPE responds more sensitively to changes in sampling steps compared to FID, further confirming that our sJPE metric effectively captures subtle variations in motion quality.
Figure 5: Time spent decoding a batch of token sequences with a batch size of 32. All tests are conducted on the same NVIDIA RTX 4090Ti. Each experiment was repeated 20 times across the entire HumanML3D test set, and the average values were reported.
@@
To evaluate the contribution of each component to model performance, we conducted ablation studies on our model’s configuration. Table 5 presents the results on the HumanML3D test set, assessing the impact of each component on reconstruction and generation quality. First, we examined alternative projection mechanisms. Replacing our projection method with standard up-convolution or repeat followed by a linear layer resulted in significant performance degradation, underscoring the effectiveness of our approach. Additionally, incorporating attention into the U-Net backbone and using full motion sequences instead of windowed motion segments resulted in performace degradation. This indicates that focusing on localized motion segments enhances the model’s generalization capability, particularly in stage 2.
@@
MoMask ${0.019}^{\pm{.001}}$ 29.5 ${0.051}^{\pm{.002}}$ ${2.957}^{\pm{.008}}$
+ DisCoRD (Ours) ${0.011}^{\pm{.000}}$ 33.3 $\mathbf{{0.032}}^{\pm{.002}}$ $\mathbf{{2.938}}^{\pm{.010}}$
Ours (Upconv) ${0.010}^{\pm{.000}}$ 31.5 ${0.039}^{\pm{.003}}$ ${2.943}^{\pm{.006}}$
Ours (Repeat & Linear) ${0.011}^{\pm{.001}}$ 31.8 ${0.038}^{\pm{.001}}$ ${2.947}^{\pm{.008}}$
Ours (w/ Attention) ${0.020}^{\pm{.000}}$ 32.7 ${0.043}^{\pm{.002}}$ ${2.983}^{\pm{.009}}$
Ours (w/o WM) $\mathbf{{0.008}}^{\pm{.000}}$ 33.4 ${0.038}^{\pm{.002}}$ ${2.952}^{\pm{.009}}$
Table 5: Ablation studies. We perform ablations on the projection, attention, and motion windowing strategies during training on the HumanML3D dataset. (WM stands for Windowed Motion.)

### v3

@@ 4 Experiments
In this section, we evaluate the effectiveness of DisCoRD in achieving motion naturalness compared to other discrete methods. We begin by assessing the naturalness of reconstructed motions to highlight the expressive capabilities of our rectified flow decoder. Then, we examine how this naturalness carries over to stage 2, generating natural motions while preserving faithfulness. We focus on text-to-motion generation due to its complex motions and diversity, but also evaluate our approach on other motion generation tasks, including co-speech gesture generation and music-to-dance generation, demonstrating the flexibility of our method.
@@ 4.1 Dataset and Evaluation
Datasets. For text-to-motion, we use HumanML3D [11] and KIT-ML [42]. HumanML3D is a 3D motion dataset with language annotations, including 14,616 motion sequences paired with 44,970 text descriptions. Sourced from motion capture data, motions are standardized to a template, scaled to 20 FPS, and cropped to 10 seconds if longer. KIT-ML is a smaller dataset with 3,911 motion sequences paired with 6,278 text descriptions. Motion capture data are downsampled to 12.5 FPS, with 1–4 descriptions per sequence. For co-speech gesture generation, we utilize the SHOW [62] dataset, while a mixed version of AIST++ [25] and HumanML3D is used for music-to-dance generation. Further dataset details are provided in the Supplementary Section B.
Figure 3: sJPE and FID response to frame-wise gaussian noise. We introduce Gaussian noise with varying standard deviations (x-axis) to ground-truth motion data and evaluate its effect on sJPE and FID. Noise sJPE is highly sensitive to subtle frame-wise perturbations, whereas Static sJPE remains low. FID is highly insensitive to frame-wise noise. Note that FID scale (y-axis, right) is very small compared to sJPE scale (y-axis, left).
Figure 4: Under-reconstruction and frame-wise noise. We visualize fine-grained motion trajectories (top), and corresponding jerk graphs (bottom), where blue and red regions indicate noise and static sJPE, respectively. Compared to other methods, DisCoRD significantly reduces sJPE, resulting in smoother motion (fewer blue regions) and greater dynamism (fewer red regions), as highlighted in green boxes.
@@
We evaluate DisCoRD on both motion reconstruction and motion generation separately. For motion reconstruction, the primary objective is to assess how effectively the decoder reconstructs motion from tokens. This is measured by Fréchet Inception Distance (FID), which assesses motion realism by comparing the feature distributions of generated and ground truth motions, and Mean Per Joint Position Error (MPJPE), which quantifies positional accuracy. For text-to-motion generation, we follow [51] and employ several established metrics: FID, R-Precision, Multimodal Distance (MM-Dist), and Multimodality (MModality). For co-speech gesture generation, we employ Fréchet Gesture Distance (FGD) [33], and for music-to-dance generation, following [53], we utilize Distk and Distg to assess the quality of generated motions by comparing the distributional spread of generated and real motions. Additional specifics on these metrics are provided in the Supplementary Section B.
Symmetric Jerk Percentage Error.
Prior works [60, 12] rely on MPJPE and FID at stage 1. However, MPJPE has limited correlation with human perceptual preferences [9], while FID, being a model-level metric extracted from pretrained network features, fails to capture per-sample naturalness [57]. Our experiments further indicate that FID is particularly insensitive to subtle, fine-grained noise (see Figure 3), which critically affects immersive motion quality [58]. To overcome these limitations, we introduce the Symmetric Jerk Percentage Error (SJPE), a metric explicitly designed to assess both under-reconstructed motions and frame-level noise through jerk. Jerk, defined as the third derivative of position with respect to time, has proven to effectively quantify subtle deviations [4, 7] and kinetic inconsistencies in motion [3, 23], being a critical measure for detecting unnatural artifacts in generated motion.
Let $J_{\text{pred},t}$ and $J_{\text{true},t}$ denote the predicted and ground truth jerk, respectively, at time $t$ over $n$ time points. Then sJPE, capturing the symmetric mean absolute percentage error [34] between predicted and ground truth jerk values, is defined as
@@
Within sJPE, we identify two components: Noise sJPE and Static sJPE. Noise sJPE corresponds to instances where $J_{\text{pred},t}>J_{\text{true},t}$, indicating an overestimation of jerk in the predicted motion, which reflects the presence of fine-grained noise. This effect is evident in discrete-based methods, where discrete tokens introduce frame-wise noise, as shown in Figure 4. Static sJPE captures instances where $J_{\text{pred},t}\leq J_{\text{true},t}$, indicating underestimation of jerk, or insufficiently dynamic predicted motion, highlighted by green boxes in Figure 4. Together, these components provide a comprehensive measure of prediction accuracy, capturing both over- and underestimations of jerk within a unified score. Additional details and results are in Supplementary Section C.
Dataset Methods 000 FID $\downarrow$ 000 MPJPE $\downarrow$ 000sJPE $\downarrow$000
H-ML3D MLD [60] (cont.) 0.017 14.7 0.404
T2M-GPT [64] 0.089 60.0 0.564
+DisCoRD(Ours) 0.031(+65%) 71.5 0.488(+13%)
MMM [41] 0.097 46.9 0.517
+DisCoRD(Ours) 0.020(+79%) 56.8 0.429(+17%)
MoMask [12] 0.019 29.5 0.512
+DisCoRD(Ours) 0.011(+42%) 33.3 0.385(+25%)
KIT-ML T2M-GPT [64] 0.470 46.4 0.526
+DisCoRD(Ours) 0.284(+40%) 58.7 0.395(+25%)
MoMask [12] 0.113 37.5 0.384
+DisCoRD(Ours) 0.103(+9%) 33.0 0.359(+7%)
Table 1: Quantitative results on motion reconstruction. DisCoRD enhances naturalness as a decoder for discrete models, shown by improvements over base models on FID and sJPE (blue). H-ML3D stands for HumanML3D and cont. for continuous.
@@ 4.2 Quantitative results
+ DisCoRD (Ours) ${0.476}^{\pm{.008}}$ ${0.663}^{\pm{.006}}$ ${0.760}^{\pm{.007}}$ ${0.095}^{\pm{.011}}$(+18%) ${3.121}^{\pm{.009}}$ ${1.831}^{\pm{.048}}$
BAMM [40] $\mathbf{{0.525}}^{\pm{.002}}$ $\mathbf{{0.720}}^{\pm{.003}}$ $\mathbf{{0.814}}^{\pm{.003}}$ ${0.055}^{\pm{.002}}$ $\mathbf{{2.919}}^{\pm{.008}}$ ${1.687}^{\pm{.051}}$
+ DisCoRD (Ours) ${0.522}^{\pm{.003}}$ $\underline{{0.715}}^{\pm{.005}}$ $\underline{{0.811}}^{\pm{.004}}$ $\underline{{0.041}}^{\pm{.002}}$ (+25%) $\underline{{2.921}}^{\pm{.015}}$ ${1.772}^{\pm{.067}}$
MoMask [12] ${0.521}^{\pm{.002}}$ ${0.713}^{\pm{.002}}$ ${0.807}^{\pm{.002}}$ ${0.045}^{\pm{.002}}$ ${2.958}^{\pm{.008}}$ ${1.241}^{\pm{.040}}$
+ DisCoRD (Ours) $\underline{{0.524}}^{\pm{.003}}$ $\underline{{0.715}}^{\pm{.003}}$ ${0.809}^{\pm{.002}}$ $\mathbf{{0.032}}^{\pm{.002}}$ (+29%) ${2.938}^{\pm{.010}}$ ${1.288}^{\pm{.043}}$
@@
+ DisCoRD (Ours) ${0.382}^{\pm{.007}}$ ${0.590}^{\pm{.007}}$ ${0.715}^{\pm{.004}}$ ${0.541}^{\pm{.038}}$ (+25%) ${3.260}^{\pm{.028}}$ $\underline{{1.928}}^{\pm{.059}}$
@@
+ DisCoRD (Ours) $\mathbf{{0.434}}^{\pm{.007}}$ $\mathbf{{0.657}}^{\pm{.005}}$ $\underline{{0.775}}^{\pm{.004}}$ $\underline{{0.169}}^{\pm{.010}}$ (+17%) $\underline{{2.792}}^{\pm{.015}}$ ${1.266}^{\pm{.046}}$
Table 2: Quantitative results on motion generation. $\pm$ indicates a 95% confidence interval. +DisCoRD indicates that the baseline model’s decoder is replaced with DisCoRD. Bold indicates the best result, while underscore refers the second best. DisCoRD improves naturalness, as evidenced by FID improvements shown in blue, while preserving faithfulness, demonstrated by R-Precision.
@@
We evaluate DisCoRD’s effectiveness in reconstructing natural motions from discrete models. Existing discrete methods often struggle to generate natural motions, as indicated by higher sJPE and FID values. While MoMask achieves competitive FID, its high sJPE suggests significant frame-wise noise and poor reconstruction quality compared to continuous models like MLD. Our proposed DisCoRD decoder substantially improves motion quality, as shown by reduced FID and sJPE metrics, overcoming the typical limitations of discrete models and producing smoother, more natural motions. Note that MPJPE measures only positional accuracy, and does not reflect motion naturalness or align with human perception [18].
@@
Table 3: Quantitative results on each method’s SHOW test set. DisCoRD outperforms baseline models on sJPE and FGD.
@@
Table 4: Quantitative results on the AIST++ test set. DisCoRD outperforms baseline model on sJPE, $\mathrm{Dist_{k}}$ and $\mathrm{Dist_{g}}$.
@@
To evaluate DisCoRD’s effectiveness in decoding predicted tokens, we use models trained in stage 1 to assess their performance in decoding tokens generated by pretrained token predictors. As shown in Table 2, our method consistently outperforms baseline models, particularly in terms of FID, achieving state-of-the-art performance in naturalness. We observe that for T2M-GPT, which employs a vanilla VQ-VAE with limited representational capacity, there is a slight decline in faithfulness, as a single token can map to multiple motions in DisCoRD. However, with RQVAEs [24], a popular quantization method on recent works [40, 12], DisCoRD performs on par and even increase faithfullness, shown by R-Precision and MM-Dist. These results indicate that when paired with a decent tokenizer, DisCoRD can significantly boost naturalness without sacrificing faithfulness, highlighting its potential as a default decoder replacement for discrete motion generation models.
@@
To validate our approach as a general method for enhancing naturalness in discrete-based human motion generation, we train DisCoRD on co-speech gesture and music-driven dance generation, conducting a comparative analysis against baseline models. As shown in Table 3 and Table 4, our method consistently outperforms baseline models across both tasks, achieving superior performance on sJPE and standard evaluation metrics. We present additional evaluation results in Supplementary Section D.
Figure 5: Decoding efficiency comparison. We report the average decoding time for a batch of 32 token sequences on an NVIDIA RTX 4090 Ti, averaged over 20 trials on the HumanML3D test set. DisCoRD achieves more better performance on motion naturalness at a comparable decoding speed to MoMask and can even decode significantly faster while maintaining superior performance.
@@
By selecting the rectified flow algorithm from among the various diffusion model variants, we exploit its efficient transport mechanism to achieve inference speeds comparable to baseline models. As shown in Figure 5, we evaluated the decoding times for both Momask and our model on tokens generated by a pretrained token generator. At the default setting of 16 sampling steps, our model achieves decoding speeds on par with MoMask while delivering superior sJPE, stage 1 FID, and stage 2 FID. Furthermore, by reducing the sampling steps, our method can decode tokens significantly fasterer than MoMask, maintaining comparable or enhanced FID and sJPE performance. Additionally, although it was not the primary focus of this experiment, we observed that sJPE responds more sensitively to changes in sampling steps compared to FID, further confirming that our sJPE metric effectively captures subtle variations in motion quality.
@@
We conduct ablation studies on DisCoRD’s each component shown in Table 5. First, we compare our decoding strategy with post refinement methods which refine output motions from MoMask’s decoder. Feedforward convolution layers show little improvement and rectified flow post refinement falls short in all metrics compared to ours. Second, we examine alternative projection mechanisms. While up-convolution or repeat followed by a linear layer show strong reconstruction performance, they fail to decode natural motion from generated token sequences (unseen at training) shown by low FID. Additionally, incorporating attention into the U-Net backbone and using full motion sequences instead of windowed motion segments result in performace degradation. This indicates that focusing on localized motion segments enhances the model’s generalization capability, particularly in stage 2.
@@
MoMask ${0.019}^{\pm{.001}}$ 0.512 ${0.051}^{\pm{.002}}$ ${2.957}^{\pm{.008}}$
+ Post Refinement (FF model) ${0.028}^{\pm{.000}}$ 0.481 ${0.044}^{\pm{.002}}$ ${2.962}^{\pm{.006}}$
+ Post Refinement (RF model) ${0.013}^{\pm{.000}}$ 0.489 ${0.035}^{\pm{.002}}$ ${2.955}^{\pm{.008}}$
+ DisCoRD (Ours) ${0.011}^{\pm{.000}}$ 0.385 $\mathbf{{0.032}}^{\pm{.002}}$ $\mathbf{{2.938}}^{\pm{.010}}$
Ours (Upconv) ${0.010}^{\pm{.000}}$ 0.375 ${0.039}^{\pm{.003}}$ ${2.943}^{\pm{.006}}$
Ours (Repeat & Linear) ${0.011}^{\pm{.001}}$ 0.342 ${0.038}^{\pm{.001}}$ ${2.947}^{\pm{.008}}$
Ours (w/ Attention) ${0.020}^{\pm{.000}}$ 0.384 ${0.043}^{\pm{.002}}$ ${2.983}^{\pm{.009}}$
Ours (w/ Full Motion Sequence) $\mathbf{{0.008}}^{\pm{.000}}$ 0.385 ${0.038}^{\pm{.002}}$ ${2.952}^{\pm{.009}}$
Table 5: Ablation studies. Evaluation on the HumanML3D test set assessing the impact of decoding strategies, projection methods, and training strategies. FF and RF denote feedforward and rectified flow model, respectively.
4.3 Qualitative Results
User Studies.
We conduct two user studies to (1) validate our motivation and method effectiveness and (2) evaluate how well sJPE aligns with human perception. The first study, shown in Figure 7, indicates that the discrete model Momask outperforms the continuous model MDM in faithfulness but lags in naturalness. In contrast, DisCoRD surpasses both, demonstrating its ability to generate motion that is both natural and faithful. In the second study, we find that sJPE exhibits 2.7 times higher correlation with human preference for naturalness compared to MPJPE, highlighting its effectiveness in evaluating sample-wise motion naturalness. Details of user studies are in Supplementary Section C.4 and E.2.
Visualizations.
We visualize motion trajectories in Figure 4, where DisCoRD, unlike discrete methods, produces smooth and expressive motion with low sJPE. Additionally, as shown in Figure 6, our method generates motions samples that closely align with textual descriptions while preserving a high degree of naturalness. Extensive visualizations are in Supplementary Section C.3 and E.1.